# Neuro-Symbolic Generation of Explanations for Robot Policies with Weighted Signal Temporal Logic

### v2

@@ Abstract
Learning-based policies have demonstrated success in many robotic applications, but often lack explainability, which poses challenges in safety-critical deployments. To address this, we propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and loose, which limits meaningful insights into the underlying policy. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three novel explainability evaluation metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification metrics. Our method is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing classification accuracy. This work bridges policy learning with explainability through formal methods, contributing to safer and more transparent decision-making in robotics.
@@ I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have significantly improved both high-level decision-making and low-level control in robotics. However, these neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. To address these concerns, various approaches have been explored to enhance explainability of learned policies [1, 2], including post-hoc analysis techniques [3, 4] and methods to learn policies that are inherently explainable [5, 6].
@@
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [7] and feature attribution [8], provide valuable information, they often lack language-based explanations, which limits human-interpretability. Autoregressive language models have recently shown promising results for providing language-based explanations [9, 10, 11], but they do not support formal verification of the output explanation, which limits their applicability in domains requiring strong safety and correctness assurances [12].
@@
Formal methods offer an alternative that emphasizes verifiability. In particular, temporal logic has been used to specify robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [13, 14, 15]. Recent efforts have also used temporal logic, primarily linear temporal logic (LTL) and signal temporal logic (STL), to infer explanations of given trajectory data. For example, [16, 17, 18] use a decision tree classifier to infer the LTL or STL specification that best explains the difference between two classes of trajectories. [19] uses a greedy explanation search method over explanation candidates in LTL driven by policy similarity metrics. However, these methods are limited in their ability to infer specifications where certain predicates are more critical than others. For instance, [18] can infer such relations only for conjunctions. Weighted signal temporal logic (wSTL) [20] extends LTL and STL by introducing importance weights over predicates, which can highlight the relative priority of various task components. The expressiveness of wSTL thus makes it particularly suitable for describing robotic behaviors in settings with competing or hierarchical objectives.
Figure 1: Existing methods for neuro-symbolic post-hoc explainability produce specifications that are verbose (i.e., include unnecessary predicates), inconsistent across training seeds, and loose (e.g., “catch-all” specifications that do not provide useful insights). Our proposed network, TLNet, addresses these concerns.
@@
The presence of real-valued weights in wSTL has enabled the use of neural network approaches for inferring wSTL specifications that explain given trajectory data. These methods have shown promising results, where they show high classification performance for various applications, including anomaly detection in naval surveillance [21], fault diagnosis [22], and weather forecasting [23]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explainability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications that do not fully capture the task or constraints). Some works address verbosity by incorporating heuristic simplification methods, like only keeping the top-$k$ weighted predicates [21, 24] or using a greedy process to remove weights [25]. However, as we show in our empirical results, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
@@
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). Our method targets a subclass of wSTL structures with arbitrary conjunctions and disjunctions of predicates in temporal clauses. TLNet includes a multi-step simplification process that reduces specification verbosity and improves consistency, while ensuring strictness is maintained. This process includes a predicate filtering step, a regularized loss function, and iterative weight pruning. We also provide theoretical justification for our wSTL-based neural network architecture, which essentially reduces learning parameters by the distributivity of importance weight, and introduce novel metrics for quantifying explainability—conciseness, consistency, and strictness—that complement standard classification accuracy, addressing a significant gap in current evaluation practices. We demonstrate our method in three simulated robotics environments that vary in continuity, dimensionality, task structure, and policy complexity, and show that it generally outperforms baselines with respect to classification accuracy and our proposed explainability metrics. TLNet bridges the gap between neural network-based policy learning and provides formal explanations of learned policies, offering a step toward safer and more transparent robotic systems.
@@ II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{>0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
@@ II-A Weighted Signal Temporal Logic
We consider wSTL syntax [20] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a function $f:S\rightarrow\mathbb{R}$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
@@ II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi_{i},$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$. We allow the weights to be zero because the zero weights do not influence the overall robustness [21].
@@
$r(\psi,\tau,t)$ ::=:= $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s_{t})-c}&\text{otherwise}\end{cases},$ (1)
$r(\neg\phi,\tau,t)$ ::=:= $-r(\phi,\tau,t),$ (2)
$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ ::=:= $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ ::=:= $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\otimes^{\mathcal{F}}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively. For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
@@
$\otimes^{\mathcal{G}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
We use $\sigma=0.5$ in this work.
@@ III-A Neural Network Architecture
ϕ=italic-$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
@@
where $w_{i}^{\mathcal{F}},w_{ij}^{\mathcal{F}},w_{i}^{\mathcal{G}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates. We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k=1:\lvert I\rvert}$ for trajectory $\tau$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [26, 19].
@@
9:     Re-normalize $\mathbf{w}$.
10:     Optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
11:     if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$/$\mathbf{w}^{\mathcal{G}}$ then break.
12:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 12.
We reduce the number of parameters that are optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
[…]
@@
where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{G}}$. We justify this weight distribution operation in Section III-E. An example TLNet architecture for a dataset containing two predicates is shown in Figure 3.
Figure 3: Example architecture of the proposed neural network when there are two predicates to consider ($\psi_{1}$ and $\psi_{2}$). $\mathbf{w}^{\mathcal{F}}$ and $\mathbf{w}^{\mathcal{G}}$ are weight vectors that are optimized. $\phi$ is the inferred explanation to the given trajectory.

### v3

@@ Abstract
Learning-based policies have demonstrated success in many robotic applications, but often lack explainability. We propose a neuro-symbolic explanation framework that generates a weighted signal temporal logic (wSTL) specification which describes a robot policy in a human-interpretable form. Existing methods typically produce explanations that are verbose and inconsistent, which hinders explainability, and are loose, which limits meaningful insights. We address these issues by introducing a simplification process consisting of predicate filtering, regularization, and iterative pruning. We also introduce three explainability metrics—conciseness, consistency, and strictness—to assess explanation quality beyond conventional classification accuracy. Our method—TLNet—is validated in three simulated robotic environments, where it outperforms baselines in generating concise, consistent, and strict wSTL explanations without sacrificing accuracy. This work bridges policy learning and explainability through formal methods, contributing to more transparent decision-making in robotics.
@@ I Introduction
Recent innovations in learning-based methods, particularly those using neural networks, have improved high-level decision-making and low-level control in robotics. However, neural network policies are inherently black-box in nature, which poses significant concerns for safety-critical robotic applications, such as autonomous vehicles, where explainability is paramount. Various approaches have been explored to enhance the explainability of learned policies [18], including post-hoc techniques [12] and methods to learn policies that are inherently explainable [14].
@@
In this paper, we focus on post-hoc analysis to enable explanation of any given robot policy. While existing approaches, like saliency maps [19], feature attribution [20], and non-temporal concept/rule extraction [23], provide valuable information, they often lack language-based explanations, which limits human interpretability. Autoregressive language models have shown promising results for providing language-based explanations [26], but they do not support formal verification of the output explanation, which limits their applicability in safety-critical domains [7].
@@
Temporal logic offers an alternative approach to specifying robotic tasks and system constraints in an interpretable yet mathematically precise manner that supports verification [8, 4, 5]. Recent efforts have also used temporal logic to infer explanations of given trajectory data. For example, [10, 6] use a decision tree (DT) classifier to infer the linear temporal logic (LTL) or signal temporal logic (STL) specification that best explains the difference between two classes of trajectories. [29] uses a greedy search over candidate LTL explanations driven by policy similarity metrics. However, these methods are limited in their ability to infer specifications where certain predicates are more critical than others. Weighted STL (wSTL) [17] extends STL by introducing importance weights over predicates, which can highlight the relative priority of various task components. The expressiveness of wSTL thus makes it particularly suitable for describing robotic behaviors in settings with competing or hierarchical objectives.
Figure 1: Existing neuro-symbolic methods often produce explanations that are verbose, inconsistent, and loose. Our method, TLNet, addresses this by generating concise, consistent, and strict wSTL specifications.
@@
The presence of real-valued weights in wSTL has enabled the use of neural networks for inferring wSTL specifications that explain given trajectory data. These methods have shown high classification accuracy for various applications, including anomaly detection [27], fault diagnosis [25], and weather forecasting [3]. However, existing work only focuses on classification accuracy, which does not fully capture the actual explainability of learned specifications. For example, as shown in Figure 1, these methods often produce specifications that are unnecessarily verbose (i.e., they include many unnecessary predicates), inconsistent across seeds, and/or loose, such that they are satisfied but do not provide useful insights (e.g., “catch-all” specifications that include many conjunctions or overly simplified specifications). Some works address verbosity by incorporating heuristic simplifications, like only keeping the top-$k$ predicates [27, 28] or using a greedy process to remove weights [15]. However, no existing methods effectively balance classification accuracy with verbosity, inconsistency, and looseness.
@@
In this work, we propose TLNet, a neural network architecture for inferring wSTL explanations of robot policies that are concise (i.e., include as few predicates as needed), consistent (i.e., produce similar explanations across training seeds), and strict (i.e., contain as few disjunctions as needed). TLNet achieves this through a multi-step simplification process consisting of predicate filtering, a regularized loss function, and iterative weight pruning. We provide theoretical justification for our network’s weight distributivity and introduce explainability metrics—conciseness, consistency, and strictness—to complement standard classification accuracy. We demonstrate our method in three simulated robotics environments, showing it outperforms baselines in our proposed metrics while matching or improving classification accuracy.
@@ II Preliminaries
Let a trajectory (or a signal) of $n$-dimensional states be a discrete-time series $\tau=s_{0},s_{1},\dots,s_{H}$, where $s_{t}\in S$ is a state at timestep $t\in\mathbb{Z}_{\geq 0}$, $S\subseteq\mathbb{R}^{n}$ is the state space, and $H$ is the time horizon of $\tau$. We also define an interval $I$ as $I:=[a,b]=\{k\in\mathbb{Z}_{\geq 0}|a\leq k\leq b;a,b\in\mathbb{Z}_{\geq 0}\}$, $\lvert I\rvert$ as the cardinality of $I$, and $t+I$ as the interval $[t+a,t+b]$.
@@ II-A Weighted Signal Temporal Logic
We consider wSTL syntax [17] defined recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi$, where $\top$ is the logical True value, $\psi:=f(s)\geq c$ is an atomic predicate given a function $f:S\rightarrow\mathbb{R}$ with constant $c\in\mathbb{R}$, and $\neg$, $\wedge$, and $\vee$ are Boolean negation, conjunction, and disjunction operators, respectively. $\mathcal{G}_{I}$ and $\mathcal{F}_{I}$ are globally and eventually operators, respectively. For conjunctive and disjunctive operators, we define the weights as $w:=[w_{i}]_{i=1:N}\in\mathbb{R}^{N}_{>0}$, where $N$ is the number of subformulae in the operator. For temporal operators, we define the weights as $w:=[w_{k}]_{k\in I}\in\mathbb{R}^{\lvert I\rvert}_{>0}$ for interval $I$. For notational simplicity, we use the following notation interchangeably: $\bigwedge_{i=1}^{N}{}^{w}\phi_{i}=\bigwedge_{i=1}^{N}w_{i}\phi_{i}$ and $\bigvee_{i=1}^{N}{}^{w}\phi_{i}=\bigvee_{i=1}^{N}w_{i}\phi_{i}$.
@@ II-B Normalized wSTL
Our neural network architecture proposed in Section III-A requires the weights within each operator to be normalized. We discuss the reasons for this requirement in Section III-E. Given this requirement, we define normalized wSTL syntax recursively as $\phi:=\top\mid\psi\mid\neg\phi\mid\bigwedge_{i=1}^{N}{}^{w}\phi_{i}\mid\bigvee_{i=1}^{N}{}^{w}\phi_{i}\mid\mathcal{G}_{I}^{w}\phi\mid\mathcal{F}_{I}^{w}\phi,$ where all terminology carries over from wSTL, except we further require $\sum_{i=1}^{N}w_{i}=\sum_{k\in I}w_{k}=1$, where $w_{i},w_{k}\in[0,1]$. We allow the weights to be zero because the zero weights do not influence the overall robustness [27].
@@
$r(\psi,\tau,t)$ $:=$ $\begin{cases}\frac{f(s_{t})-c}{\sup_{s\in S}f(s)-c}&\text{if $f(s_{t})-c\geq 0$,}\\ \frac{f(s_{t})-c}{\inf_{s\in S}f(s)-c}&\text{otherwise}\end{cases},$ (1)
$r(\neg\phi,\tau,t)$ $:=$ $-r(\phi,\tau,t),$ (2)
$r\left(\bigwedge_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ $:=$ $\otimes^{\wedge}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (3)
$r\left(\bigvee_{i=1}^{N}{}^{w}\phi_{i},\tau,t\right)$ $:=$ $\oplus^{\vee}(w,[r(\phi_{i},\tau,t)]_{i=1:N}),$ (4)
$r\left(\mathcal{G}^{w}_{I}\phi,\tau,t\right)$ $:=$ $\otimes^{\mathcal{G}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (5)
$r\left(\mathcal{F}^{w}_{I}\phi,\tau,t\right)$ $:=$ $\oplus^{\mathcal{F}}_{I}(w,[r(\phi,\tau,t^{\prime})]_{t^{\prime}\in t+I}),$ (6)
where the aggregation functions $\otimes^{\wedge}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\oplus^{\vee}:\mathbb{R}^{N}_{\geq 0}\times\mathbb{R}^{N}\rightarrow\mathbb{R},\otimes^{\mathcal{G}}_{I}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$, and $\oplus^{\mathcal{F}}_{I}:\mathbb{R}^{\lvert I\rvert}_{\geq 0}\times\mathbb{R}^{\lvert I\rvert}\rightarrow\mathbb{R}$ correspond to the operators $\wedge,\vee,\mathcal{G}$, and $\mathcal{F}$, respectively. In practice, the infimum and supremum bounds in Equation 1 can be inferred from physical properties of the environment (we assume known map sizes in our results) or sampled trajectories. For notational simplicity, we interchangeably use $r(\phi,\tau,t)$ and $r_{\tau t}^{\phi}$.
@@
$\otimes^{\mathcal{G}}_{I}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{-r_{\tau t^{\prime}}^{\phi}}{\sigma})},$ (9)
$\oplus^{\mathcal{F}}_{I}(w,[r_{\tau t^{\prime}}^{\phi}]_{t^{\prime}\in t+I})$ $:=\frac{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}r_{\tau t^{\prime}}^{\phi}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}{\sum_{t^{\prime}\in t+I}w_{t^{\prime}}\exp(\frac{r_{\tau t^{\prime}}^{\phi}}{\sigma})}.$ (10)
We empirically found that decreasing $\sigma$ better approximates the max/min aggregations, but can increase training instability due to larger gradients, while increasing $\sigma$ too much can lead to nonsensical explanations.
@@ III-A Neural Network Architecture
$\phi=$ $0.5\mathcal{F}_{I}^{\bar{w}}\left[\bigwedge_{i=1}^{N_{\mathrm{AP}}}w_{i}^{\mathcal{F}}\bigvee_{j=1}^{2N_{\mathrm{AP}}}w_{ij}^{\mathcal{F}}\psi_{j}\right]$
@@
where $w_{i}^{\mathcal{F}},w_{ij}^{\mathcal{F}},w_{i}^{\mathcal{G}},w_{ij}^{\mathcal{G}}\in[0,1]$ are importance weights and $\psi_{j}\in\Psi\cup\Psi_{\neg}$ are predefined atomic predicates. We define the weights $\bar{w}$ for temporal operators over interval $I=[0,H]$ to be $\bar{w}=[\bar{w}_{k}=1/\lvert I\rvert]_{k\in I}$. We define the conjunction weights of the temporal clauses to be 0.5 to equally weigh the importance of the task ($\mathcal{F}$) and constraint ($\mathcal{G}$) parts of the explanation. We choose this class of structures because they allow us to represent a general class of robot policies aiming to complete a task while satisfying a set of constraints [16, 29].
@@
9:  Re-normalize and optimize $\mathbf{w}$ to minimize $\tilde{\mathcal{L}}$.
10:  if only one non-zero weight in $\mathbf{w}^{\mathcal{F}}$$\mathbf{w}^{\mathcal{G}}$ then break.
11:return $\phi$ generated from $\mathbf{w}$ and $\Psi$ using Equation 12.
We reduce the number of parameters to be optimized by distributing the weights in Equation 11, resulting in a specification of the following form,
$\phi_{\mathbf{w}}=$
@@
where $\mathbf{w}=\mathbf{w}^{\mathcal{F}}\mathbin{\|}\mathbf{w}^{\mathcal{G}}$, $\mathbf{w}^{\mathcal{F}}=[\tilde{w}_{ij}^{\mathcal{F}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\mathbf{w}^{\mathcal{G}}=[\tilde{w}_{ij}^{\mathcal{G}}]_{i=1:N_{\mathrm{AP}},j=1:2N_{\mathrm{AP}}}$, $\tilde{w}_{ij}^{\mathcal{F}}=w_{i}^{\mathcal{F}}w_{ij}^{\mathcal{F}}$, and $\tilde{w}_{ij}^{\mathcal{G}}=w_{i}^{\mathcal{G}}w_{ij}^{\mathcal{G}}$. This weight distribution reduces the number of optimized parameters from $2N_{\mathrm{AP}}+4N_{\mathrm{AP}}^{2}$ to $4N_{\mathrm{AP}}^{2}$. We justify this weight distribution operation in Section III-E. Figure 3 shows a TLNet architecture for a dataset containing two predicates.
Figure 3: Example TLNet architecture with two predicates, where $\phi_{\mathbf{w}}$ is the explanation generated from $\mathbf{w}$.